Vishaka and Vaishali are shopping for antique jewellery. Vishaka’s 

demand function for antique jewellery is Q = 30 - 3P, and Vaishali’s demand 

function is Q = 48 - 4P. What will their combined market demand be if the 

price is Rs 5?


1
Expert's answer
2021-05-19T20:29:31-0400

"Q = 30 - 3P"

"Q = 30 - 3(5)"

"Q = 30 - 15"

"Q = 15"


"Q = 48 - 4P"

"Q = 48 - 4(5)"

"Q = 48 - 20"

"Q = 28"


combined demand will be

"15+28=43"